Solve the following equations:
(p+1)(2p-3)+2=0

Answer by checkley79(3341)


(p+1)(2p-3)+2=0
2P^2+2P-3P-3+2=0
2P^2-P-1=0
(2P+1)(P-1)=0 ANS.